the sum of two numbers is 42 the greater number is three more than twice the lesser number. what are the numbers?
constructing equations
how do i construct this to a equation?

Answer by stanbon(75887)

the sum of two numbers is 42
the greater number is three more than twice the lesser number.
what are the numbers?
constructing equations
how do i construct this to a equation?
---------------
Let one of the numbers be "x".
Then the other number is "42-x".
----
Equation:
42-x = 2x + 3
3x = 39
x = 13 (one of the numbers)
42-x = 29 (the other number)
===============================
Cheers,
Stan H.